OpenStudy (anonymous):

In the coming year, a vehicle manufacturer has decided to manufacture 150 vehicles per day. The function v = 150d represents the company’s production for the coming year, v, with respect to the number of days, d. The rate of change of the function representing the number of vehicles manufactured for the coming year is ? and its graph is a ?. So, the function is a ? function.

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (michele_laino):

Hint: I think that we have to compute the number \(N\) of manufactured vehicles inside one year, so we have: \(N=150 \times 365=...?\)

OpenStudy (anonymous):

ok i got it

OpenStudy (michele_laino):

Since the manufacturing company works for, I think, \(315\) days per year, then better is if we do this computation: \(N=150×315=...?\), since on Saturdays and on Sundays, the manufactueing company doesn't work

OpenStudy (michele_laino):

so, please what is \(150 \times 315=...?\)

OpenStudy (anonymous):

47250

OpenStudy (michele_laino):

Sorry, again, since we the manufacturing compnay is working for 5 days per week and one year contains 52 weeks, then the manufacturing company works \(260\) days in total. So in one year, such manufacturing company produces \(150 \times 260=39,000\) vehicles

OpenStudy (michele_laino):

so the the requested production is \(39,000\) vehicles per year